OpenStudy (anonymous):

1. Find the length of segment AB if the coordinates at A are (6,1) and B are (-2,3). Give your answer in simplest radical form.

OpenStudy (anonymous):

pythagoras for this one. over 8, up 2 get \[\sqrt{8^2+2^2}\]

OpenStudy (anonymous):

what? i got the square root of -60 and idk how to simplify

OpenStudy (anonymous):

\[\sqrt{64+4}=\sqrt{68}\]

OpenStudy (anonymous):

you can write \[2\sqrt{17}\] if you like